Summary for dayshift 9-9-99

Richard_Coleman (coleman@fsgi03)
Thu, 09 Sep 1999 15:24:24 CDT

Shift Summary for day shift, 9-9-99
On Shift: Arun,RickC

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0.75 hr laser scan
Accounting of lost time:

========================================================================
Runs ended this shift: 14474,14475(laser scan)

Tapes brought to FCC this shift:online split tapes
QKCR07,QKLA16,QKMU13,QKPE14,QKSM17,QKA023-026,QKNL31-33,QKN033-036,
QKCC18-19

========================================================================
Accesses during shift: none

========================================================================
Overview of shift: Laser Scan, cosmics, very quiet

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14474 7:00 Beam shutdown begins

14475 7:20 Laser scan

7:22 Ran AN832NEGTOPOS to prepare for mu run
following shutdown

13:10 Vaulted split tapes

13:25 Paged Dave N. about West side Ar/Et empty
He will stop by later.

14:30 Asked MCR if shutdown was ending on schedule
at 4:00 pm... too soon to tell- many keys still out
14:45 x-view fault is in the readout again, seems fine
after 3 hrs of cosmics-Jason.

======================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift